# Fully Synthetic Data for Complex Surveys

### v3

@@ 1 Introduction
Researchers have proposed a variety of approaches for generating fully synthetic data in complex surveys. The suggestion in early work (Rubin, 1993; Raghunathan et al., 2003; Reiter, 2002) was to take a Bayesian finite population inference approach, in which the agency (1) builds predictive models for the survey variables conditional on design features like stratum/cluster indicators or size measures, which are assumed known by the agency for every unit in the population, (2) imputes the missing survey variables for the nonsampled units in the population, and (3) takes a simple random sample from the completed population to release as one synthetic data set. A related approach uses the weighted finite population Bayesian bootstrap (WFPBB) (Dong et al., 2014), in which the agency generates completed populations by replicating individuals from the confidential data in proportion to their survey weights and then releases the completed populations, forgoing the step of simple random sampling. More recently, it has been suggested to build synthetic data models that account for the sampling design directly, so that they estimate the joint distribution of the population data. For example, the agency can use a pseudo-likelihood approach (Pfeffermann, 1993; Savitsky and Toth, 2016), in which each individual’s contribution to the likelihood function of a synthesis model is raised to a power that is a function of the survey weights (Kim et al., 2021). Departing from the proposal of Rubin (1993), a completely different approach is to create and attach new weights to synthetic data records simulated from models that are agnostic to the survey weights (United Nations Economic Commission for Europe, 2022). Here, the goal is to allow users to use weighted estimates that scale up to the finite population. The new weights can be created by treating the survey weights as a variable in the synthesis, so that the agency specifies a predictive model for the weights. The simulated weights may be adjusted by raking or calibration before inclusion in the released file.
@@ 2.2.2 Derivation with SynRep-1
The pairs of equations (2.2.25) and (2.2.9), (2.2.26) and (2.2.10), and (2.2.27) and 2.2.13 can be viewed as equivalent when $R=1$.
@@ 3 Simulation Studies
In this section, we present simulation studies to illustrate the repeated sampling properties of the inferential methods in Section 2.2.1 and 2.2.2 for various finite population quantities.
@@ 3.1 Simulation design
From this population, we sample $\mathcal{D}$ using a PPS sample of size $n=500$ survey units, setting $\pi_{i}=nx_{i}/\sum_{i=1}^{N}x_{i}$ and using the function “ppss” in the $R$ package “pps” (Gambino, 2021). Under this PPS sampling design, we expect that unweighted inferences using $\mathcal{D}$ should be badly biased for $(\bar{Y}_{1},\bar{Y}_{2})$ but perhaps not so for $\beta$. We repeat the sampling process to create 500 independent realizations of $\mathcal{D}$.
@@
For each $\mathcal{D}$, we implement SynRep-R and SynRep-1 with various $(M,R)$. Specifically, we examine $(M=4,R=5)$, $(M=10,R=5)$, $(M=50,R=5)$, $(M=10,R=10)$, $(M=10,R=25)$, and $(M=10,R=50)$. Of course, the choice of $R$ only affects SynRep-R. We implement the WFPBB using the “polyapost” package in $R$ (Meeden et al., 2020), creating pseudo-populations $(\mathcal{P}^{(1)},\dots,\mathcal{P}^{(M)})$ each comprising 25,000 individuals. From each $\mathcal{P}^{(m)}$ where $m=1,\dots,M$, we take a simple random sample of size $n$ to make a corresponding $\mathcal{D}^{(m)}$. To make each synthetic data replicate stemming from each $\mathcal{D}^{(m)}$, we sample $n$ synthetic values for $Y_{1}$ using a Bernoulli distribution with probability set to the empirical proportion of $Y_{1}$ in $\mathcal{D}^{(m)}$. We sample the corresponding synthetic values of $Y_{2}$ from normal distributions with means equal to the predicted values from the regression of $Y_{2}$ on $Y_{1}$, computed using the synthetic values of $Y_{1}$ and the unbiased estimates of the coefficients computed with $\mathcal{D}^{(m)}$, and variance equal to the unbiased estimate of the regression variance computed with $\mathcal{D}^{(m)}$.
@@
To assist in evaluating the repeated sampling performances of SynRep-1 and SynRep-R, we also use results computed with $\mathcal{P}_{pseudo}$ and $\mathcal{D}_{srs}$. Specifically, in each of the 500 simulation runs, we define Pseudo-Pop as the procedure that uses a point estimator of $\bar{Q}$ and variance estimator of $(1+1/M)B$ computed with the WFPBB-generated pseudo-populations $(\mathcal{P}^{(1)},\dots,\mathcal{P}^{(M)})$. We define Pseudo-SRS as the procedure that uses a point estimator of $\bar{q}$ and variance estimator of Raghunathan et al. (2003) computed with $(\mathcal{D}^{(1)},\dots,\mathcal{D}^{(M)})$. As a comparison against what happens if we disregard the sampling design entirely, we define SRSsyn as the procedure that generates synthetic data by using (i) the unweighted sample proportion for $Y_{1}$ as the Bernoulli probability to generate $n$ synthetic values of $Y_{1}$ and (ii) the unweighted estimates of parameters in the regression of $Y_{2}$ on $Y_{1}$ as the parameters of the normal distribution to generate the corresponding $n$ synthetic values of $Y_{2}$.
@@
We also evaluate the repeated sampling performances of pseudo-likelihood approaches to making fully synthetic data. For each synthesis model, i.e., the Bernoulli and linear regression models, we start with a likelihood function defined as the product of the contributions from each individual in $\mathcal{D}$. We create the pseudo-likelihood by raising each individual’s contribution to a power defined by the individual’s survey weight. We use these weighted pseudo-likelihoods to estimate synthesis model parameters. We implement this approach using the software $Stan$ (Stan Development Team, 2024), which can generate posterior samples of model parameters based on user-specified likelihood functions. We run $Stan$ to create four chains of 4,000 iterations and discard the first 2,000 iterations as burn-in. We randomly sample one of the resulting draws and use its parameter values in the Bernoulli and linear regression models to generate the synthetic data. We repeat this process $M$ times and apply the inference rules in Raghunathan et al. (2003). We call this method Wtreg.
@@
Let superscript $s$ index the results from simulation run $s$, where $s=1,\dots,500$. For any estimator $\hat{q}$ for any of the methods we examine, we compute the percent bias, $100\sum_{s=1}^{500}(\hat{q}^{s}-Q)/(500Q)$. We compute the proportion of the 500 95% confidence intervals based on $\hat{q}$ and its corresponding variance estimate that cover $Q$. We also compute the ratio of the empirical variance of the 500 values of $\hat{q}$ to the empirical variance of the 500 values of the HT point estimator. To investigate the accuracy of variance estimators, for each method we compute the ratio of the average of the 500 variance estimates over its corresponding empirical variance. Finally, to examine the stability of the variance estimator for each method, we compute the standard deviation of the 500 variance estimates. We present results for the first four quantities in the main text and for the last quantity in the Appendix.
@@ 3.2 Results
We first investigate the properties of SynRep-R and SynRep-1 for the various settings of $(M,R)$. Figure 3.1, Figure 3.2, and Figure 3.3 display results for $\bar{Y}_{1}$, $\bar{Y}_{2}$, and $\beta$, respectively, for these two methods as well as for Pseudo-Pop and Pseudo-SRS. All four methods offer approximately unbiased point estimates of the three finite population quantities, with simulated percent biases all around 1% or lower. These small biases originate primarily from the step of completing populations, as the biases in Pseudo-Pop are close to the biases in the other three methods. As expected, compared to the variance for HT, the simulated variances are increasingly inflated as $M$ decreases. Holding $M=10$ constant, decreasing $R$ tends to increase the simulated variances, although the effects are less pronounced than those from decreasing $M$. Taken together, these results suggest it is preferable to increase $M$ rather than $R$ when keeping $MR$ constant. For example, when we compare SynRep-R with $(M=10,R=5)$ to SynRep-1 with $M=50$, the latter tends to result in smaller empirical variance with closer-to-nominal coverage rates. Similar benefits appear when comparing SynRep-R with $(M=10,R=25)$ to SynRep-R with $(M=50,R=5)$. This finding accords with results from Reiter (2008), who considered a similar trade-off for nested multiple imputation for partially synthetic and missing data. We note that using larger values of $M$ also offers smaller variability in the estimated variances, as shown in the Appendix.
@@
$(M,R)$ Method $\bar{X}$ $\bar{Y}$ $\beta$
M4R5 Pseudo-SRS 0.09 0.16 0.14
M4R5 SynRep-R 0.13 0.14 0.13
M4R5 SynRep-1 0.20 0.25 0.23
M10R5 Pseudo-SRS 0.01 0.02 0.03
M10R5 SynRep-R 0.02 0.02 0.02
M10R5 SynRep-1 0.04 0.07 0.06
@@
We next turn to compare SynRep-R and SynRep-1 with other approaches, particularly Wtreg, Wtreg-Boot, and SRSsyn. Here, we set $M=10$ and, where relevant, $R=10$. Figure 3.4 summarizes the repeated sampling performances of the methods that account for survey weights. For all these methods, the point estimators have simulated percent biases that typically are not more than 1%. For SynRep-R and SynRep-1, the average variance estimates are close to their corresponding empirical variances, and the coverage rates are close to nominal. For Wtreg and Wtreg-Boot, the variance estimators can underestimate the corresponding empirical variances severely, especially for $\bar{Y}_{1}$ and $\bar{Y}_{2}$, resulting in confidence interval coverage rates that can be substantially lower than the nominal 95% level. The bootstrap step in Wtreg-Boot results in more reliable variance estimates compared to Wtreg, but Wtreg-Boot is not as well calibrated as SynRep-R and SynRep-1, which have closer to nominal coverage rates. As expected, HT results in accurate estimates with near nominal coverage rates. We note that Figure 3.4 does not display results for Direct and SRSsyn because they perform poorly for $\bar{Y}_{1}$ and $\bar{Y}_{2}$. For these two methods, the simulated biases for $\bar{Y}_{1}$ and for $\bar{Y}_{2}$ are around $16$% and $11$%, respectively, with coverage rates near 0 and near 30%, respectively. These results emphasize the importance of accounting for informative designs when generating fully synthetic data that can be analyzed as simple random samples.
@@ 5 Discussion
Fourth, negative values of the variance estimates can arise, particularly for small $M$, resulting in the need for ad hoc adjustments. Future work can investigate inferential methods based on sampling from the distributions used in the derivations, as illustrated for fully synthetic data by  Si and Reiter (2011).
@@ Appendix A Appendix: Additional Simulation Results
Figure A.1 displays the variability of the 500 values of estimated variances of the point estimators for $\beta,\bar{Y}_{1}$, and $\bar{Y}_{2}$ for the simulation with the PPS design. The variability tends to decrease with $M$. Increasing $R$ when $M$ is held constant seems not to have much impact on the stability of the results. We see increased variability as the procedures introduce more steps that involve randomness; that is, as we go from Pseudo-Pop to Pseudo-SRS to SynRep-R and SynRep-1. The variability tends to be largest for SynRep-1.

### v4

@@ 1 Introduction
Researchers have proposed a variety of approaches for generating fully synthetic data in complex surveys. The suggestion in early work (Rubin, 1993; Raghunathan et al., 2003; Reiter, 2002) was to take a Bayesian finite population inference approach, in which the agency (i) builds predictive models for the survey variables conditional on design features like stratum/cluster indicators or size measures, which are assumed known by the agency for every unit in the population, (ii) imputes the missing survey variables for the nonsampled units in the population, and (iii) takes a simple random sample from the completed population to release as one synthetic data set. A related approach uses the weighted finite population Bayesian bootstrap (WFPBB) (Dong et al., 2014), in which the agency generates completed populations by replicating individuals from the confidential data in proportion to their survey weights and then releases the completed populations, forgoing the step of simple random sampling. More recently, it has been suggested to build synthetic data models that account for the sampling design directly, so that they estimate the joint distribution of the population data. For example, the agency can use a pseudo-likelihood approach (Pfeffermann, 1993; Savitsky and Toth, 2016), in which each individual’s contribution to the likelihood function of a synthesis model is raised to a power that is a function of the survey weights (Kim et al., 2021). Departing from the proposal of Rubin (1993), a completely different approach is to create and attach new weights to synthetic data records simulated from models that are agnostic to the survey weights (United Nations Economic Commission for Europe, 2022). Here, the goal is to allow users to use weighted estimates that scale up to the finite population. The new weights can be created by treating the survey weights as a variable in the synthesis, so that the agency specifies a predictive model for the weights. The simulated weights may be adjusted by raking or calibration before inclusion in the released file.
@@ 2.2.2 Derivation with SynRep-1
The pairs of equations (2.2.25) and (2.2.9), (2.2.26) and (2.2.10), and (2.2.27) and (2.2.13) can be viewed as equivalent when $R=1$.
@@ 3 Simulation Studies
In this section, we present simulation studies to illustrate the repeated sampling properties of the inferential methods in Section 2.2 for various finite population quantities.
@@ 3.1 Simulation design
From this population, we sample $\mathcal{D}$ using a PPS sample of size $n=500$ survey units, setting $\pi_{i}=nx_{i}/\sum_{i=1}^{N}x_{i}$ and using the function “ppss” in the $R$ package “pps” (Gambino, 2021). Under this PPS sampling design, we expect that unweighted inferences using $\mathcal{D}$ should be badly biased for $(\bar{Y}_{1},\bar{Y}_{2})$ but perhaps not so for $\beta$. We repeat the sampling process to create 1000 independent realizations of $\mathcal{D}$.
@@
For each $\mathcal{D}$, we implement SynRep-R and SynRep-1 with various $(M,R)$. Specifically, we examine $(M=4,R=5)$, $(M=10,R=5)$, $(M=50,R=5)$, $(M=10,R=10)$, $(M=10,R=25)$, and $(M=10,R=50)$. The choice of $R$ only affects SynRep-R. We implement the WFPBB using the “polyapost” package in $R$ (Meeden et al., 2020), creating pseudo-populations $(\mathcal{P}^{(1)},\dots,\mathcal{P}^{(M)})$ each comprising 25,000 individuals. From each $\mathcal{P}^{(m)}$ where $m=1,\dots,M$, we take a simple random sample of size $n$ to make a corresponding $\mathcal{D}^{(m)}$. To make each synthetic data replicate stemming from each $\mathcal{D}^{(m)}$, we sample $n$ synthetic values for $Y_{1}$ using a Bernoulli distribution with probability set to the empirical proportion of $Y_{1}$ in $\mathcal{D}^{(m)}$. We sample the corresponding synthetic values of $Y_{2}$ from normal distributions with means equal to the predicted values from the regression of $Y_{2}$ on $Y_{1}$, computed using the synthetic values of $Y_{1}$ and the unbiased estimates of the coefficients computed with $\mathcal{D}^{(m)}$, and variance equal to the unbiased estimate of the regression variance computed with $\mathcal{D}^{(m)}$.
@@
To assist in evaluating the repeated sampling performances of SynRep-1 and SynRep-R, we also use results computed with $\mathcal{P}_{pseudo}$ and $\mathcal{D}_{srs}$. Specifically, in each of the 1000 simulation runs, we define Pseudo-Pop as the procedure that uses a point estimator of $\bar{Q}$ and variance estimator of $(1+1/M)B$ computed with the WFPBB-generated pseudo-populations $(\mathcal{P}^{(1)},\dots,\mathcal{P}^{(M)})$. We define Pseudo-SRS as the procedure that uses a point estimator of $\bar{q}$ and variance estimator of Raghunathan et al. (2003) computed with $(\mathcal{D}^{(1)},\dots,\mathcal{D}^{(M)})$. As a comparison against what happens if we disregard the sampling design entirely, we define SRSsyn as the procedure that generates synthetic data by using (i) the unweighted sample proportion for $Y_{1}$ as the Bernoulli probability to generate $n$ synthetic values of $Y_{1}$ and (ii) the unweighted estimates of parameters in the regression of $Y_{2}$ on $Y_{1}$ as the parameters of the normal distribution to generate the corresponding $n$ synthetic values of $Y_{2}$.
@@
We also evaluate the repeated sampling performances of pseudo-likelihood approaches to making fully synthetic data. For each synthesis model, i.e., the Bernoulli and linear regression models, we start with a likelihood function defined as the product of the contributions from each individual in $\mathcal{D}$. We create the pseudo-likelihood by raising each individual’s contribution to a power defined by the individual’s survey weight. We use these weighted pseudo-likelihoods to estimate synthesis model parameters. We implement this approach using the software $Stan$ (Stan Development Team, 2024), which can generate posterior samples of model parameters based on user-specified likelihood functions. We run $Stan$ to create four chains of 4,000 iterations and discard the first 2,000 iterations as burn-in. We randomly sample one of the resulting draws and use its parameter values in the Bernoulli and linear regression models to generate the synthetic data. We repeat this process $M$ times and apply the inference rules in Raghunathan et al. (2003). We call this method Wtreg. We note that that Kim et al. (2021) use the variance estimator (2.2) from Raghunathan et al. (2003) with $\bar{v}=0$. Kim et al. (2021) release synthetic populations (where $\bar{v}=0$) rather than synthetic samples (where $\bar{v}>0$).
@@
Let superscript $s$ index the results from simulation run $s$, where $s=1,\dots,1000$. For any estimator $\hat{q}$ for any of the methods we examine, we compute the percent bias, $100\sum_{s=1}^{1000}(\hat{q}^{s}-Q)/(1000Q)$. We compute the proportion of the 1000 95% confidence intervals based on $\hat{q}$ and its corresponding variance estimate that cover $Q$. We also compute the ratio of the empirical variance of the 1000 values of $\hat{q}$ to the empirical variance of the 1000 values of the HT point estimator. To investigate the accuracy of variance estimators, for each method we compute the ratio of the average of the 1000 variance estimates over its corresponding empirical variance. Finally, to examine the stability of the variance estimator for each method, we compute the standard deviation of the 1000 variance estimates. We present results for the first four quantities in the main text and for the last quantity in the Appendix.
@@ 3.2 Results
We first investigate the properties of SynRep-R and SynRep-1 for the various settings of $(M,R)$. Figure 3.1, Figure 3.2, and Figure 3.3 display results for $\bar{Y}_{1}$, $\bar{Y}_{2}$, and $\beta$, respectively, for these two methods as well as for Pseudo-Pop and Pseudo-SRS. All four methods offer approximately unbiased point estimates of the three finite population quantities, with simulated percent biases generally around 1% or lower. These small biases originate primarily from the step of completing populations, as the biases in Pseudo-Pop are close to the biases in the other three methods. As expected, compared to the variance for HT, the simulated variances are increasingly inflated as $M$ decreases. Holding $M=10$ constant, decreasing $R$ tends to increase the simulated variances, although the effects are less pronounced than those from decreasing $M$. The variability in SynRep-1 results with fixed $M$ reflects Monte Carlo error. Taken together, these results suggest it is preferable to increase $M$ rather than $R$ when keeping $MR$ constant. For example, when we compare SynRep-R with $(M=10,R=5)$ to SynRep-1 with $M=50$, the latter tends to result in smaller empirical variance with closer-to-nominal coverage rates. Similar benefits appear when comparing SynRep-R with $(M=10,R=25)$ to SynRep-R with $(M=50,R=5)$. This finding accords with results from Reiter (2008), who considered a similar trade-off for nested multiple imputation for partially synthetic and missing data. We note that using larger values of $M$ also offers smaller variability in the estimated variances, as shown in the Appendix.
@@
$(M,R)$ Method $\bar{Y}_{1}$ $\bar{Y}_{2}$ $\beta$
M4R5 Pseudo-SRS 0.09 0.15 0.13
M4R5 SynRep-R 0.11 0.17 0.13
M4R5 SynRep-1 0.17 0.26 0.22
M10R5 Pseudo-SRS 0.01 0.02 0.02
M10R5 SynRep-R 0.01 0.03 0.02
M10R5 SynRep-1 0.04 0.09 0.07
@@
We next turn to compare SynRep-R and SynRep-1 with other approaches, particularly Wtreg, Wtreg-Boot, and SRSsyn. Here, we set $M=10$ and, where relevant, $R=10$, and draw 500 repeated samples. Figure 3.4 summarizes the repeated sampling performances of the methods that account for survey weights. For all these methods, the point estimators have simulated percent biases that typically are negligible. For SynRep-R and SynRep-1, the average variance estimates are close to their corresponding empirical variances, and the coverage rates are close to nominal. For Wtreg and Wtreg-Boot, the variance estimators can underestimate the corresponding empirical variances severely, especially for $\bar{Y}_{1}$ and $\bar{Y}_{2}$, resulting in confidence interval coverage rates that can be substantially lower than the nominal 95% level. The bootstrap step in Wtreg-Boot results in more reliable variance estimates compared to Wtreg, but Wtreg-Boot is not as well calibrated as SynRep-R and SynRep-1, which have closer to nominal coverage rates. As expected, HT results in accurate estimates with near nominal coverage rates. We note that Figure 3.4 does not display results for Direct and SRSsyn because they perform poorly for $\bar{Y}_{1}$ and $\bar{Y}_{2}$. For these two methods, the simulated biases for $\bar{Y}_{1}$ and for $\bar{Y}_{2}$ are around $16$% and $11$%, respectively, with coverage rates near 0 and near 30%, respectively. These results emphasize the importance of accounting for informative designs when generating fully synthetic data that can be analyzed as simple random samples.
@@ 5 Discussion
Fourth, we present ad hoc adjustments to deal with negative values of the variance estimates. We may be able to improve on those adjustments. For example, we may be able to adapt the strategy in Si and Reiter (2011), who develop inferential methods for fully synthetic data based on sampling from the distributions used in the derivations of the combining rules. Additionally, as pointed out by a reviewer, it may be beneficial to use the insight of Raab et al. (2018) for the sampling and synthesis components of the derivation in SynRep-R. This results in an alternative variance estimator, $\left(1+M^{-1}\right)b_{syn}-\left(1+R^{-1}\right)\bar{v}_{syn}.$ Future work can investigate the performance of these alternative inference methods.
@@ Appendix A Appendix: Additional Simulation Results
Figure A.1 displays the variability of the 1000 values of estimated variances of the point estimators for $\beta,\bar{Y}_{1}$, and $\bar{Y}_{2}$ for the simulation with the PPS design. The variability tends to decrease with $M$. Increasing $R$ when $M$ is held constant seems not to have much impact on the stability of the results. We see increased variability as the procedures introduce more steps that involve randomness; that is, as we go from Pseudo-Pop to Pseudo-SRS to SynRep-R and SynRep-1. The variability tends to be largest for SynRep-1.